The longest side of a triangle is 11 inches, the second side is 8 inches shorter than that. Find the 3rd side.
Answer by nerdybill(5404)

The longest side of a triangle is 11 inches, the second side is 8 inches shorter than that. Find the 3rd side.
.
You must apply Pythagorean theorem.
The longest side (11 inches) is the hypothenuse.
The second side (since it is 8 inches shorter) is 3 inches.
.
Therefore:
Let x = third side
.
Then we have:
x^2 + 3^2 = 11^2
x^2 + 9 = 121
x^2 = 112
x = {-10.583, 10.583}
.
Tossing out the negative answer:
3rd side is 10.583 inches